OpenStudy (anonymous):

Given the equation sqrt(3)/xy = 2/(x^2-y^2) change the equation to polar form and simplify the answer by using fundamental identities. Where do I begin?

OpenStudy (anonymous):

Do I change it as:\[\frac{ \sqrt{3} }{ rcos \theta rsin \theta }=\frac{ 2 }{ (rcos \theta)^{2}-(rsin \theta)^{2}}\]

OpenStudy (anonymous):

Any luck @rorytsm?

OpenStudy (anonymous):

@RadEn, perhaps you may know?

OpenStudy (anonymous):

i know you can distribute the exponent on the right side of the equation and then factor out an r^2 or -r^2 and use the identity that cos^2theta+sin^2theta = 1 i'm just not 100% sure about which to factor out to get rid of the - between the two so i decided not to post

OpenStudy (anonymous):

you can also factor an r out on the left hand side for a basic simplification, but other than that i'm not really too sure what else you can do

OpenStudy (anonymous):

Well you had the same idea as I did. I'm not sure how to change the sqrt(3) and 2 to polar

OpenStudy (anonymous):

pretty sure those don't change, in all the problems using spherical and cylindrical coordinates that i've worked so far this semester, we never changed any "normal" numbers over to polar coordinates, they were always just left as is.

OpenStudy (anonymous):

unless you want to divide each side by two to get sqrt(3)/2 on the left which is equal to sin(pi/3)

OpenStudy (anonymous):

so then you'd have (sin(pi))/3r(costhetasintheta)

OpenStudy (anonymous):

I could make the left side:\[\frac{ \sqrt(3) }{ r(\frac{ 1 }{ 2 }\sin2 \theta )}\]correct?

OpenStudy (anonymous):

i'm pretty sure you can

OpenStudy (anonymous):

So then I can make the left side \[\frac{ 2 }{ r(\cos2 \theta) }\]Correct?

OpenStudy (anonymous):

I mean right side

OpenStudy (anonymous):

your r would have to be squared, and i really don't think there's many other options other than looking for the cos^2+sin^2 identity there unless you went with two half angle formulas (1-cos2theta)/2 and (1+cos2theta)/2 and then combined those

OpenStudy (anonymous):

the cos2theta's would cancel leaving you with 2/2 = 1 but either way you get the same result